# Tidal perturbations of an extreme mass ratio inspiral around a Kerr black hole

### v1

@@ 6.2 LR shifts
By solving Eqs. (6.19) to zeroth order in the tidal parameter $\eta$, we recover the known results for the radial coordinate, impact parameter, and frequency of the unperturbed LRs of a Kerr black hole [Chandrasekhar:1985kt, Bardeen:1972fi]
@@
Note that, in the non-spinning limit $a=0$, this expression correctly reduces to the first-order tidal correction to the light-ring frequency of a Schwarzschild black hole derived in Ref. [Camilloni:2023rra].
The analytical results obtained for the tidal deformations of the ISCO and the LR are derived for a generic external tidal environment. The only assumptions are that the characteristic length scale of the Kerr black hole, set by its mass $M$, is much smaller than the curvature radius $\mathscr{R}$ of the external spacetime ($M\ll\mathscr{R}$), and that the tidal perturbation is retained only up to the leading quadrupolar order.
@@
This behavior is consistent across all orbital quantities considered for both the ISCO and the LR, and admits a clear physical interpretation. As originally discussed by Bardeen and collaborators [Bardeen:1972fi], increasing the black-hole spin causes prograde circular orbits to move closer to the event horizon $r_{+}$, while retrograde orbits are pushed to larger radii. As a consequence, prograde ISCO and LR orbits probe regions of stronger gravitational binding, where the coupling to a fixed external tidal field is effectively reduced, leading to a suppression of tidal effects as $a$ increases. Conversely, retrograde orbits reside at larger radii for higher spins, where the tidal interaction is stronger, resulting in enhanced tidal deformations.
@@ 7 Conclusions
In this paper, we developed an analytic framework to describe the tidal deformation of rotating black holes and applied it to the study of the secular dynamics of test particles in the strong-field regime. By solving the Teukolsky master equation in the Hartle–Hawking tetrad for static modes and employing metric reconstruction techniques in the outgoing radiation gauge, we derived fully explicit expressions for the quadrupolar metric perturbation of a Kerr black hole subject to slowly varying external tidal fields. In the non-spinning limit, our results smoothly reduce to the known expressions for tidally deformed Schwarzschild black holes [Binnington:2009bb]. All results are obtained in the near-zone regime $M\ll\mathscr{R}$ and at leading quadrupolar order in the tidal deformation.
Using the reconstructed metric, we investigated the secular dynamics of test particles in the tidally deformed Kerr spacetime and derived the corresponding secular Hamiltonian to first order in the tidal perturbation. This allowed us to compute tidal-induced shifts in the location and properties of the innermost stable circular orbit and the light ring. We found that these shifts exhibit a strong dependence on the black-hole spin and on the relative orientation between the orbital angular momentum and the spin. In particular, tidal effects are suppressed for prograde orbits around rapidly rotating black holes, while they are significantly enhanced for retrograde orbits. These results highlight the role of spin-tidal couplings in shaping the strong-field orbital dynamics of matter and radiation in the near-horizon region, and extend to spinning black holes the strong-field tidal effects previously identified in charged black-hole spacetimes [Grilli:2024fds].
@@
Several directions for future work emerge from this analysis. An important extension is the inclusion of higher-order tidal multipoles, as well as the study of time-dependent tidal perturbations. It would also be relevant to compute the associated energy and angular-momentum fluxes, following existing approaches in black-hole perturbation theory [Poisson:2004cw, Chatziioannou:2012gq, Berens:2024czo]. These extensions are likely to be particularly relevant for the modeling of EMRIs and for future tests of general relativity with space-based gravitational-wave detectors such as LISA.
@@
A related but distinct application concerns hierarchical triple systems, in which the external tidal field is generated by a third massive black hole. Recent work has shown that tidal interactions in such systems can significantly affect the eccentricity evolution and merger timescale of compact binaries treated within post-Newtonian theory [Camilloni:2023xvf, Cocco:2025adu, Cocco:2025udb]. The tidally deformed Kerr metric derived here provides a strong-field framework to study analogous effects when the inner object is modeled as a test particle orbiting a spinning black hole, thereby enabling a systematic comparison between EMRI and comparable-mass regimes. We plan to explore these extensions in future work.
@@ Acknowledgements
M. Cocco, G. Grignani, and M. Orselli acknowledge financial support from the Italian Ministry of University and Research (MUR) through the program “Dipartimenti di Eccellenza 2018–2022” (Grant SUPER-C). G. Grignani and M. Orselli also acknowledge support from the “Fondo di Ricerca d’Ateneo” 2023 (GraMB) of the University of Perugia, and from the Italian Ministry of University and Research (MUR) via PRIN 2022ZHYFA2, GRavitational wavEform models for coalescing compAct binaries with eccenTricity (GREAT).
M. Cocco, T. Harmark, M. van de Meent, and M. Orselli acknowledge support from the “Center of Gravity”, a Center of Excellence funded by the Danish National Research Foundation under Grant No. DNRF184. M. van de Meent further acknowledges financial support from the VILLUM Foundation (Grant No. VIL37766), the DNRF Chair program (Grant No. DNRF162) funded by the Danish National Research Foundation, and the European Union’s Horizon ERC Synergy Grant Making Sense of the Unexpected in the Gravitational-Wave Sky (GWSky–101167314).

### v2

@@ 6.2 LR shifts
By solving Eqs. (6.19) to zeroth order in the tidal parameter $\eta$, we recover the known results for the radial coordinate, impact parameter, and frequency of the unperturbed LRs of a Kerr black hole [5, 75]
@@
Note that, in the non-spinning limit $a=0$, this expression correctly reduces to the first-order tidal correction to the light-ring frequency of a Schwarzschild black hole derived in Ref. [73].
The analytical results obtained for the tidal deformations of the ISCO and the LR are derived for a generic external tidal environment. The only assumptions are that $M\ll\mathscr{R}$, and that the tidal perturbation is retained only up to the leading quadrupolar order.
@@
This behavior is consistent across all orbital quantities considered for both the ISCO and the LR, and admits a clear physical interpretation. As originally discussed by Bardeen and collaborators [75], increasing the black-hole spin causes prograde circular orbits to move closer to the event horizon $r_{+}$, while retrograde orbits are pushed to larger radii. As a consequence, prograde ISCO and LR orbits probe regions of stronger gravitational binding, where the coupling to a fixed external tidal field is effectively reduced, leading to a suppression of tidal effects as $a$ increases. A qualitatively similar suppression of tidal effects in the extremal limit was observed in Ref. [74] for charged, non-rotating black-hole binaries. Conversely, retrograde orbits reside at larger radii for higher spins, where the tidal interaction is stronger, resulting in enhanced tidal deformations.
@@ 7 Conclusions
In this paper, we developed an analytic framework to describe the tidal deformation of rotating black holes and applied it to the study of the secular dynamics of test particles in the strong-field regime. By solving the Teukolsky master equation in the Hartle–Hawking tetrad for static modes and employing metric reconstruction techniques in the outgoing radiation gauge, we derived fully explicit expressions for the quadrupolar metric perturbation of a Kerr black hole subject to slowly varying external tidal fields. In the non-spinning limit, our results smoothly reduce to the known expressions for tidally deformed Schwarzschild black holes [23]. All results are obtained in the near-zone regime $M\ll\mathscr{R}$ and at leading quadrupolar order in the tidal deformation.
Using the reconstructed metric, we investigated the secular dynamics of test particles in the tidally deformed Kerr spacetime and derived the corresponding secular Hamiltonian to first order in the tidal perturbation. This allowed us to compute tidal-induced shifts in the location and properties of the innermost stable circular orbit and the light ring. We found that these shifts exhibit a strong dependence on the black-hole spin and on the relative orientation between the orbital angular momentum and the spin. In particular, tidal effects are suppressed for prograde orbits around rapidly rotating black holes, while they are significantly enhanced for retrograde orbits. These results highlight the role of spin-tidal couplings in shaping the strong-field orbital dynamics of matter and radiation in the near-horizon region, and extend to spinning black holes the strong-field tidal effects previously identified in charged black-hole spacetimes [74].
@@
Several directions for future work emerge from this analysis. An important extension is the inclusion of higher-order tidal multipoles, as well as the study of time-dependent tidal perturbations. It would also be relevant to compute the associated energy and angular-momentum fluxes, following existing approaches in black-hole perturbation theory [9, 13, 42]. These extensions are likely to be particularly relevant for the modeling of EMRIs and for future tests of general relativity with space-based gravitational-wave detectors such as LISA.
@@
A related but distinct application concerns hierarchical triple systems, in which the external tidal field is generated by a third massive black hole. Recent work has shown that tidal interactions in such systems can significantly affect the eccentricity evolution and merger timescale of compact binaries treated within post-Newtonian theory [94, 95, 96]. The tidally deformed Kerr metric derived here provides a strong-field framework to study analogous effects when the inner object is modeled as a test particle orbiting a spinning black hole, thereby enabling a systematic comparison between EMRI and comparable-mass regimes.
Finally, from a more fundamental perspective, it is desirable to understand the physics of near-extremal and extremal black holes. These have been shown to be very sensitive to perturbations, which can be amplified drastically close to the event horizon [97]. It would be interesting to explore if, in the extremal limit, there exist observables where tidal perturbations are similarly amplified, even within general relativity. We plan to explore these extensions in future work.
@@ Acknowledgements
M. Cocco, G. Grignani, and M. Orselli acknowledge financial support from the Italian Ministry of University and Research (MUR) through the program “Dipartimenti di Eccellenza 2018–2022” (Grant SUPER-C). G. Grignani and M. Orselli also acknowledge support from the “Fondo di Ricerca d’Ateneo” 2023 (GraMB) of the University of Perugia, and from the Italian Ministry of University and Research (MUR) via PRIN 2022ZHYFA2, GRavitational wavEform models for coalescing compAct binaries with eccenTricity (GREAT). M. Cocco, T. Harmark, M. van de Meent, and M. Orselli acknowledge support from the “Center of Gravity”, a Center of Excellence funded by the Danish National Research Foundation under Grant No. DNRF184. D. Pereñiguez is further supported by NSF Grants No. AST-2307146, PHY-2513337, PHY-090003, and PHY-20043, by NASA Grant No. 21-ATP21-0010, by John Templeton Foundation Grant No. 62840, by the Simons Foundation, and by Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. M. van de Meent further acknowledges financial support from the VILLUM Foundation (Grant No. VIL37766), the DNRF Chair program (Grant No. DNRF162) funded by the Danish National Research Foundation, and the European Union’s Horizon ERC Synergy Grant Making Sense of the Unexpected in the Gravitational-Wave Sky (GWSky–101167314).